OpenStudy (anonymous):

Solve: ----> (2-y)^3

OpenStudy (campbell_st):

it has to be equal to something to solve it

OpenStudy (anonymous):

Its to the power of 3 -_-

OpenStudy (campbell_st):

yes... but that's not an equation... an equation needs an equal sign to be able to solve it. do you need to simplify, distribute... substitute... or something else?

OpenStudy (anonymous):

factorise

OpenStudy (anonymous):

idk man its my HW lol

OpenStudy (campbell_st):

its already factorised.... a binomial of degree 3

OpenStudy (anonymous):

ya true

OpenStudy (anonymous):

in class we solved a question like this i just didnt understand

OpenStudy (campbell_st):

(2 - y)^3 = 27 is an equation you can solve

OpenStudy (anonymous):

something like (2-y)(2-y)(2-y)

OpenStudy (triciaal):

did you take any notes?

OpenStudy (anonymous):

we all can solve (2-y)^2 right

OpenStudy (anonymous):

just idk how with ^3

OpenStudy (triciaal):

maybe you need to expand and write the polynomial?

OpenStudy (triciaal):

(2-y)^3 = (2-y)^2*(2-y) what is (2-y)^2 =

OpenStudy (anonymous):

right okay we r getting somewhere

OpenStudy (anonymous):

(2-y)^2 = (2-y)(2-y) = 4-2y-2y-y^2

OpenStudy (anonymous):

=4-4y-y^2

OpenStudy (anonymous):

wat now :P

OpenStudy (triciaal):

when you get that now multiply by (2-y) remember the distributive property you just used now you have 3 terms

OpenStudy (anonymous):

and how exactly ?

OpenStudy (anonymous):

4-4y-y^2 (2-y) like so ?

OpenStudy (triciaal):

(2-y)^2 = 4 -2y + y^2 (2-y)^3 = (2-y)(2-y)^2 = (2-y)(4 - 2y + y^2) = 2(4 - 2y + y^2) -y(4 - 2y + y^2) yes

OpenStudy (triciaal):

advise to use ( ) otherwise think only -y^2 *(2-y)

OpenStudy (triciaal):

continue from where I was when I answered yes

OpenStudy (anonymous):

alright sure thanks

OpenStudy (anonymous):

so like this (2-y)(4-2y+y^2) 8-4y+2y^2-4y+2y^2-y^3 ?

OpenStudy (triciaal):

simplify by grouping like terms common practice to rearrange in order of degree (highest to lowest etc) are you back in class now

OpenStudy (anonymous):

no at home

OpenStudy (anonymous):

ok group like terms

OpenStudy (anonymous):

ok i think i got the answere

OpenStudy (anonymous):

8-8y+4y^2-Y^3 final answer ? how does it look?

OpenStudy (triciaal):

I have the same i need to check

OpenStudy (anonymous):

great, most likely correct then :D

OpenStudy (triciaal):

8 -12y + 6 y^2 + y^3

OpenStudy (triciaal):

yes found the error!!

OpenStudy (triciaal):

final correct answer is 8 -12y + 6y^2 - y^3

OpenStudy (triciaal):

@IDK69 the method is correct my error was using (4-2y + y^2) instead of (4-4y + y^2)

OpenStudy (triciaal):

I'm so sorry if I let you get it wrong. you expanded correctly. we had the same answer because you continued on my wrong entry. I knew something was wrong because of the pattern for the result.

OpenStudy (anonymous):

ohh ok thats no problem appreciate the help